A heat engine produces 30, 000 J of mechanical work per second of operation. If its efficiency is 20% how much heat is discharged from this engine per second?  


1
Expert's answer
2021-04-29T10:41:39-0400

"\eta=1-\frac{Q_2}{Q_1}\implies"

"Q_2=(1-\eta)Q_1,"

"Q_2=(1-0.2)\cdot 30000=24000~J."